Tania is as twice as old as Anna. The sum of their ages is 51. how old is Tania?
Answer by mananth(16946)

Let Anna's age = x
Tania's age = 2x
Their sum = 2x+x=3x
..
2x+x=51
3x=51
/3
3x/3=51/3
x=17 years Anna's age
Tania's age = 2x
2*17=34 years
...
m.ananth@hotmail.ca